# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ 4 Detection of coherent structures from MHD to sub-ion scales
In this section we describe the methodology to detect the structures from MHD down to sub-ion scales.
@@ 4.1 Local intermittency measure
Figure 3: A 30 minutes zoom, $[03:10,03:40]$ UT, within the analyzed time interval of 5 hours on November 6, 2018. From top to bottom: (a) magnetic field in RTN reference frame, (b) LIM of the magnetic fluctuations of the total energy $L(t,\tau)$ Equation (8), (c) artificial magnetic field $\mathbf{B}_{rand}$ with random phases and the same Fourier amplitudes as original magnetic field measurements, (d) LIM of the artificial signal $L_{rand}(t,\tau)$, (e) the comparison of the integrated LIMs $I(t)=\langle L(t,\tau)\rangle_{\tau\in[10^{-2},10^{3}]~{}s}$ (blue), and the $I_{rand}(t)=\langle L_{rand}(t,\tau)\rangle_{\tau\in[10^{-2},10^{3}]~{}s}$ (black). The horizontal red line shows $I_{threshold}=\text{max}(I_{rand}(t))$ as defined in Figure 4.
@@
We use the Local Intermittency Measure (LIM) $L(t,\tau)$ (Farge, 1992) based on Morlet wavelets in order to detect the structures. The value $L(t,\tau)$ shows the total energy of fluctuations at a given moment in time $t$ at a given time scale $\tau$, relative to the average energy at that scale:
@@
In Figure 3 we show a 30 minutes zoom within $T^{\prime}$. Panel (a) gives RTN components of the measured $\mathbf{B}$. Panel (b) shows the observed $L(t,\tau)$. The vertical elongations of enhanced $L(t,\tau)$ values are due to coupled (or coherent) phases of the fluctuations (Lion et al., 2016; Perrone et al., 2016; Alexandrova, 2020). Indeed, to see this point better, we construct an artificial signal that has the same Fourier spectrum as the original magnetic field measurements, but with random phases (Hada et al., 2003; Koga & Hada, 2003). This synthetic signal $\mathbf{B}_{rand}$ is shown in Figure 3(c), while the corresponding LIM $L_{rand}(t,\tau)$ is shown in the panel (d). The energy distribution of the synthetic signal is incoherent (randomly distributed in the $(t,\tau$)–plane), i.e., peaks of $L_{rand}(t,\tau)$ at different $\tau$ are not observed at the same time. Therefore, the vertical elongations in the observed $L(t,\tau)$ correspond to magnetic fluctuations with coupled phases across scales where the elongation is observed. The high energy of these events with respect to the mean is a sign of intense coherent structures formed in the turbulent medium (e.g. Farge, 1992; Bruno, 2019). So, we observe coherent structures which extend from inertial to sub-ion timescales. Using the Taylor hypothesis, the timescale range $\tau\in\tau_{all}=[10^{-2},10^{3}]$ s can be converted into the spatial range $\ell=V\cdot\tau\in[3,3\cdot 10^{5}]$ km, or $\ell/\rho_{i}=6[0.1,10^{4}]$ in terms of the ion Larmor radius ($\rho_{i}=5$ km).
@@
Figure 3(e) shows $I(t)$ (blue line), random phased integrated LIM $I_{rand}(t)$ (black line) and the threshold $I_{threshold}=\text{max}(I_{rand}(t))$ (red horizontal line). The local maxima of $I(t)>I_{threshold}$ give the central times of the coherent structures present in the original signal. We refer below this method as the integrated LIM selection.
The comparison of original $I(t)$ and random phased $I_{rand}(t)$ distributions is shown on Figure 4. The $I_{rand}$ distribution (in black) is close to Gaussian with a mean of 1 (because of the normalization and random phases). On the contrary, $I(t)$ (in blue-azure) has a long tail of extreme values due to the presence of coherent structures integrated over all time scales.
@@
$P=\text{Time}(I(t)>I_{threshold})/T^{\prime},$
@@
In this paper, we will also use the integrated LIM over the reduced time-scale ranges, to understand in more details the nature of the structures at MHD, ion and subion scales, where physics is different. So, we can define integrated LIM $I_{i}=(I_{MHD},I_{ion},I_{subion})$ over the corresponding range of timescales $\tau_{j}=(\tau_{MHD},\tau_{ion},\tau_{subion})$, defined in Equation (6):
@@
This band-integrated LIM selection allows us to see how the number of the structures and filling factor changes with scale band.
@@
We find a relatively small number of MHD scale structures ($N_{MHD}=$196) with high filling factor ($P_{MHD}=$12%), compared to $P_{ion}=$7% and $P_{sub-ion}=$6% for much more numerous ion scale structures ($N_{ion}=2028$) and sub-ion scale structures ($N_{sub-ion}=11167$). We remark, that our estimations of $P$ are conservative, as far as only time where LIM is over the threshold is counted, but the structure’s field decreasing from its center exist outside of the time where the energy of the structure is concentrated, e.g., (Perrone et al., 2016). So, the filling factor can be more than twice larger than given here. Finally, numerous small scale events populate larger ones and may exist outside them as well.
@@ 4.2 Magnetic field at different scales
Table 1: Main parameters of the distributions shown in Figure 4 for the magnetic field components in RTN coordinates. From left to right: standard deviation $\sigma_{RTN}$ [nT], average standard deviation $\sigma_{noise}$ [nT] (in absence of coherent structures), skewness $s_{RTN}$, average absolute skewness $\langle|s|\rangle$, kurtosis $\kappa_{RTN}$, and the average kurtosis $\langle\kappa\rangle$.
Thanks to Morlet wavelets and LIM we know now the central times of the structures covering all scales and the ones within different scale bands. In order to study magnetic field fluctuations $\delta\mathbf{B}$ in the physical space around these central times, within different scale bands, we use band-pass filter for fluctuations on frequency ranges given by Equation (6) and shown by color bands in Figure 2. We complete this analysis by studying the large scale fluctuations of $\mathbf{B}_{lowpass}-\mathbf{B}_{0}$ where the mean field $\mathbf{B}_{0}$ is defined as the average field over the time interval $T^{\prime}$. We use finite impulse response (FIR) Hamming low-pass filter with a cut-off frequency of $10^{-2}$ Hz to calculate the large scale magnetic field fluctuations of $\delta\mathbf{B}=\mathbf{B}_{lowpass}-\mathbf{B}_{0}$ (Smith, 1997).
Figure 5 shows distributions of the filtered magnetic field (solid lines) compared to the filtered signal with random phases (dashed lines). Panel (a) shows the lowpass-filtered fluctuations of the magnetic field. Panels (b-d) show the FIR Hamming bandpass-filtered fluctuations on MHD, ion, and sub-ion scales, respectively. At each band of scales we characterise the amplitude of incoherent fluctuations as follows:
@@
The random phase signal fluctuations have Gaussian distributions at all scales (Figure 5). The observed $\delta\mathbf{B}$ show scale-dependent deviation from Gaussianity. Table 1 gives the moments of the observed distributions for 3 components and at 4 different scale ranges. The distributions have non-zero skewness $s$ (a normalized measure of a distribution asymmetry). The fourth normalized moment, kurtosis $\kappa$, increases from 3-4 at large scales, up to 12-24 at sub-ion scales. In comparison, Gaussian noise has $s=0$ and $\kappa=3$.
Distributions of lowpass magnetic field components are asymmetric with respect to zero, especially radial and tangential (see Figure 5(a)). The skewness of those components has high absolute value and opposite signs: $s_{R}=1.1$ and $s_{T}=-1.1$. The lowpass magnetic field distributions doesn’t have pronounced non-gaussian tails, so the kurtosis $\kappa$ are slightly above 3, so close to the gaussian noise value (see Table 1)).
In the inertial and smaller scale ranges, the distributions have weaker asymmetry ($\langle|s|\rangle=(|s_{R}|+|s_{T}|+|s_{N}|)/3\leq 0.2$). Non-gaussian tails are identified at MHD scales (Figure 5(b)) and become even more pronounced at ion and sub-ion scales (Figure 5(c-d)). The kurtosis $\kappa_{T}$ and $\kappa_{N}$ monotonically increase from MHD to subion scales (see Table 1). The kurtosis of the radial magnetic field $\kappa_{R}$ component is growing from MHD to ion scales and then decrease at subion scales. This behavior of $\kappa_{R}$ can be explained by the proximity of the SCM noise, which starts to influence $\delta B_{R,subion}$, the weakest of the 3 components of magnetic fluctuations at these scales, see the red PDF in Figure 5(d).
@@ 5 Model structures
In this Section we discuss several models of the coherent structures. This gives us a necessary background to determine the dominant type of structures in the large statistics of events. These models have been developed in the MHD framework. Therefore they are not applicable a priori for ion and subion scales. But kinetic-scale turbulence may be described with fluid-like equations, which are structurally similar to reduced MHD equations. Therefore, it is reasonable to expect that similar types of structures (vortices, current sheets) can be distinguished among coherent structures on ion and subion scales. This is why we compare these models of coherent structures at kinetic scales as well.
@@ 5.1.1 Monopole Alfvén vortex
Figure 6 (column (a), top row) shows the crossing of a monopole vortex (Petviashvili & Pokhotelov, 1992), details of the model are described in appendix A.1.1. The set of trajectories, selected to cross the vortex, is shown by the blue transparent cone on the top left panel of Figure 6. The set is parametrised by the angle $\phi$. Two trajectories, central ($\phi=0^{\circ}$) and off-center ($\phi=10^{\circ}$), are shown in black and blue lines correspondingly. The second panel from the top shows the three magnetic field components of the monopolar vortex crossed by a spacecraft along the black trajectory in the top panel. The third panel shows the dependencies $B_{x}(B_{y})$ and $B_{z}(B_{y})$ for both central and off-center trajectories (black and blue lines respectively). The off-center trajectory has ’clover’-like polarisation in $B_{x}(B_{y})$ (blue curve). In case of the crossing through the center, the polarisation is linear (black line).
Figure 6 (column (a), bottom row) shows the MVA eigen-value ratio $\lambda_{3}/\lambda_{2}$ as a function of $\lambda_{2}/\lambda_{1}$ for 50 different trajectories (see the blue cone in the top panel). The eigen-values are ordered as $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$, with the eigen-vector $\mathbf{e_{3}}$ being the minimum variance direction. The color between violet and yellow indicates the angle $\phi$ of the trajectories: $\phi=0^{\circ}$ corresponds to the crossing through the center and $\phi=25^{\circ}$, to the side-crossing. In this plot we test the effect of an added noise with a relative amplitude $\epsilon$ defined by
@@
where $\delta B_{noise}$ is the noise amplitude and $\delta B_{\perp}$ is the amplitude of the vortex. The eigenvalue ratios $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$ are dependent on $\epsilon$. The results for two levels of noise are shown: $\epsilon_{1}=0.001$ with filled circles, and $\epsilon_{2}=0.1$ with crosses. In case of negligible noise, the points are located along the x-axis. For larger $\epsilon$, the eigenvalues become more comparable (they are identical for a noise dominated case), then the points move toward the upper right corner on the plane $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$. In Section 7, $\epsilon$ is estimated from observations.
@@ 5.1.2 Dipole Alfvén vortex
Figure 6 column (b), top panel shows the magnetic field of the dipole vortex. This type of Alfven vortices is particular, because it’s axis is inclined with respect to the background magnetic field and it is propagating, see details of the model in appendix A.1.2. The magnetic field components are symmetric in time around the vortex axis (while for the monopole vortex they are anti-symmetric). The magnetic polarisation (third panel from the top) is different for the crossing at the vortex center (black trajectory) and the side crossing (blue trajectory).
Figure 6 (column (b), bottom panel) gives the minimal variance eigenvalues ratios for two noise levels. In case of the low noise, $\epsilon=0.001$, $\lambda_{3}/\lambda_{2}\sim 0$ and $\lambda_{2}/\lambda_{1}\in[0.1,0.6]$ (filled circles). For $\epsilon=0.1$, as for the monopole vortex, both ratios increase: the points in the $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$–plane move towards the upper right corner. In case of random noise, $\lambda_{1}\simeq\lambda_{2}\simeq\lambda_{3}$. Thus in the $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$–plane the random noise occupies the corner $(1,1)$.

### v4

@@ 4 Detection of coherent structures from MHD to sub-ion scales
In this section, we describe the methodology to detect the structures from MHD down to sub-ion scales.
@@ 4.1 Local intermittency measure
Figure 3: A 30 minutes zoom, [03:10,03:40] UT, within the analyzed time interval of 5 hours on November 6, 2018. From top to bottom: (a) magnetic field in the RTN reference frame, (b) LIM of the magnetic fluctuations of the total energy $L(t,\tau)$, Equation (8), (c) artificial magnetic field $\mathbf{B}_{rand}$ with random phases and the same Fourier amplitudes as original magnetic field measurements, (d) LIM of the artificial signal $L_{rand}(t,\tau)$, (e) the comparison of the integrated LIMs $I(t)=\langle L(t,\tau)\rangle_{\tau\in[10^{-2},10^{3}]~{}s}$ (blue), and the $I_{rand}(t)=\langle L_{rand}(t,\tau)\rangle_{\tau\in[10^{-2},10^{3}]~{}s}$ (black). The horizontal red line shows $I_{threshold}=\text{max}(I_{rand}(t))$ as defined in Figure 4.
@@
We use the Local Intermittency Measure (LIM) $L(t,\tau)$ (Farge, 1992) based on Morlet wavelets to detect the structures. The value $L(t,\tau)$ shows the total energy of fluctuations at a given moment in time $t$ at a given time scale $\tau$, relative to the average energy at that scale:
@@
In Figure 3 we show a 30-minute zoom within $T^{\prime}$. Panel (a) gives RTN components of the measured $\mathbf{B}$. Panel (b) shows the observed $L(t,\tau)$. The vertical elongations of enhanced $L(t,\tau)$ values are due to coupled (or coherent) phases of the fluctuations (Lion et al., 2016; Perrone et al., 2016; Alexandrova, 2020). Indeed, to see this point better, we construct an artificial signal that has the same Fourier spectrum as the original magnetic field measurements, but with random phases (Hada et al., 2003; Koga & Hada, 2003). This synthetic signal $\mathbf{B}_{rand}$ is shown in Figure 3(c), while the corresponding LIM $L_{rand}(t,\tau)$ is shown in the panel (d). The energy distribution of the synthetic signal is incoherent (randomly distributed in the $(t,\tau$)–plane), i.e., peaks of $L_{rand}(t,\tau)$ at different $\tau$ are not observed at the same time. Therefore, the vertical elongations in the observed $L(t,\tau)$ correspond to magnetic fluctuations with coupled phases across scales where the elongation is observed. The high energy of these events with respect to the mean is a sign of intense coherent structures formed in the turbulent medium (e.g., Farge, 1992; Bruno, 2019). So, we observe coherent structures which extend from inertial to sub-ion timescales. Using the Taylor hypothesis, the timescale range $\tau\in\tau_{all}=[10^{-2},10^{3}]$ s can be converted into the spatial range $\ell=V\cdot\tau\in[3,3\cdot 10^{5}]$ km, or $\ell/\rho_{i}=6[0.1,10^{4}]$ in terms of the ion Larmor radius ($\rho_{i}=5$ km).
@@
Figure 3(e) shows $I(t)$ (blue line), random phased integrated LIM $I_{rand}(t)$ (black line) and the threshold $I_{threshold}=\text{max}(I_{rand}(t))$ (red horizontal line). The local maxima of $I(t)>I_{threshold}$ gives the central times of the coherent structures present in the original signal. In the following, we refer to this method as the integrated LIM selection.
The comparison of original $I(t)$ and random phased $I_{rand}(t)$ distributions is shown in Figure 4. The $I_{rand}$ distribution (in black) is close to Gaussian with a mean of 1 (because of the normalization and random phases). On the contrary, $I(t)$ (in blue-azure) has a long tail of extreme values due to the presence of coherent structures integrated over all time scales.
@@
$P=\text{Time}(I(t)>I_{threshold})/T^{\prime},$ (10)
@@
In this paper, we will also use the integrated LIM over the reduced time-scale ranges, to understand in more detail the nature of the structures at MHD, ion, and subion scales, where physics is different. So, we can define integrated LIM $I_{i}=(I_{MHD},I_{ion},I_{subion})$ over the corresponding range of timescales $\tau_{j}=(\tau_{MHD},\tau_{ion},\tau_{subion})$, defined in Equation (6):
@@
This band-integrated LIM selection allows us to see how the number of the structures and filling factor changes with the band of scales.
@@
We find a relatively small number of MHD scale structures ($N_{MHD}=$196) with high filling factor ($P_{MHD}=$12%), compared to $P_{ion}=$7% and $P_{sub-ion}=$6% for much more numerous ion scale structures ($N_{ion}=2028$) and sub-ion scale structures ($N_{sub-ion}=11167$). We remark, that our estimations of $P$ are conservative, as far as only time where LIM is over the threshold is counted, but the structure’s field decreasing from its center exists outside of the time where the energy of the structure is concentrated, e.g., (Perrone et al., 2016). So, the filling factor can be more than twice as large as indicated herein. Finally, numerous small-scale events populate larger ones and may exist outside them as well.
@@ 4.2 Magnetic field at different scales
Table 1: Main parameters of the distributions shown in Figure 4 for the magnetic field components in the RTN coordinates. From left to right: standard deviation $\sigma_{RTN}$ [nT], average standard deviation $\sigma_{noise}$ [nT] (in absence of coherent structures), skewness $s_{RTN}$, average absolute skewness $\langle|s|\rangle$, kurtosis $\kappa_{RTN}$, and the average kurtosis $\langle\kappa\rangle$.
Thanks to Morlet wavelets and LIM we know now the central times of the structures covering all scales and the ones within different scale bands. In order to study magnetic field fluctuations $\delta\mathbf{B}$ in the physical space around these central times, within different scale bands, we use a band-pass filter for fluctuations on frequency ranges given by Equation (6) and shown by color bands in Figure 2. We complete this analysis by studying the large scale fluctuations of $\mathbf{B}_{lowpass}-\mathbf{B}_{0}$ where the mean field $\mathbf{B}_{0}$ is defined as the average field over the time interval $T^{\prime}$. We use finite impulse response (FIR) Hamming low-pass filter with a cut-off frequency of $10^{-2}$ Hz to calculate the large scale magnetic field fluctuations of $\delta\mathbf{B}=\mathbf{B}_{lowpass}-\mathbf{B}_{0}$ (Smith, 1997).
Figure 5 shows distributions of the filtered magnetic field (solid lines) compared to the filtered signal with random phases (dashed lines). Panel (a) shows the lowpass-filtered fluctuations of the magnetic field. Panels (b-d) show the FIR Hamming bandpass-filtered fluctuations on MHD, ion, and sub-ion scales, respectively. At each band of scales, we characterize the amplitude of incoherent fluctuations as follows:
@@
The random phase signal fluctuations have Gaussian distributions at all scales (Figure 5). The observed $\delta\mathbf{B}$ show scale-dependent deviation from Gaussianity. Table 1 gives the moments of the observed distributions for 3 components at 4 different scale ranges. The distributions have non-zero skewness $s$ (a normalized measure of a distribution asymmetry). The fourth normalized moment, kurtosis $\kappa$, increases from 3-4 at large scales, up to 12-24 at sub-ion scales. In comparison, Gaussian noise has $s=0$ and $\kappa=3$.
Distributions of lowpass magnetic field components are asymmetric with respect to zero, especially radial and tangential (see Figure 5(a)). The skewness of those components has high absolute value and opposite signs: $s_{R}=1.1$ and $s_{T}=-1.1$. The lowpass magnetic field distributions don’t have pronounced non-gaussian tails, so the kurtosis $\kappa$ is slightly above 3, so close to the gaussian noise value (see Table 1)).
In the inertial and smaller scale ranges, the distributions have weaker asymmetry ($\langle|s|\rangle=(|s_{R}|+|s_{T}|+|s_{N}|)/3\leq 0.2$). Non-gaussian tails are identified at MHD scales (Figure 5(b)) and become even more pronounced at ion and sub-ion scales (Figure 5(c-d)). The kurtosis $\kappa_{T}$ and $\kappa_{N}$ monotonically increase from MHD to subion scales (see Table 1). The kurtosis of the radial magnetic field $\kappa_{R}$ component is growing from MHD to ion scales and then decreases at subion scales. This behavior of $\kappa_{R}$ can be explained by the proximity of the SCM noise, which starts to influence $\delta B_{R,subion}$, the weakest of the 3 components of magnetic fluctuations at these scales, see the red PDF in Figure 5(d).
@@ 5 Model structures
In this Section, we discuss several models of the coherent structures. This gives us a necessary background to determine the dominant type of structures in the large statistics of events. These models have been developed in the MHD framework. Therefore they are not applicable a priori for ion and subion scales. But kinetic-scale turbulence may be described with fluid-like equations, which are structurally similar to reduced MHD equations. Therefore, it is reasonable to expect that similar types of structures (vortices, current sheets) can be distinguished among coherent structures on ion and subion scales. This is why we compare these models of coherent structures at kinetic scales as well.
@@ 5.1.1 Monopole Alfvén vortex
Figure 6 (column (a), top row) shows the crossing of a monopole vortex (Petviashvili & Pokhotelov, 1992), details of the model are described in appendix A.1.1. The set of trajectories, selected to cross the vortex, is shown by the blue transparent cone on the top left panel of Figure 6. The set is parameterized by the angle $\phi$. Two trajectories, central ($\phi=0^{\circ}$) and off-center ($\phi=10^{\circ}$), are shown in black and blue lines correspondingly. The second panel from the top shows the three magnetic field components of the monopolar vortex crossed by a spacecraft along the black trajectory in the top panel. The third panel shows the dependencies $B_{x}(B_{y})$ and $B_{z}(B_{y})$ for both central and off-center trajectories (black and blue lines respectively). The off-center trajectory has ’clover’-like polarisation in $B_{x}(B_{y})$ (blue curve). In the case of the crossing through the center, the polarisation is linear (black line).
Figure 6 (column (a), bottom row) shows the MVA eigenvalue ratio $\lambda_{3}/\lambda_{2}$ as a function of $\lambda_{2}/\lambda_{1}$ for 50 different trajectories (see the blue cone in the top panel). The eigenvalues are ordered as $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$, with the eigenvector $\mathbf{e_{3}}$ being the minimum variance direction. The color between violet and yellow indicates the angle $\phi$ of the trajectories: $\phi=0^{\circ}$ corresponds to the crossing through the center and $\phi=25^{\circ}$, to the side-crossing. In this plot, we test the effect of an added noise with a relative amplitude $\epsilon$ defined by
@@
where $\delta B_{noise}$ is the noise amplitude and $\delta B_{\perp}$ is the amplitude of the vortex. The eigenvalue ratios $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$ are dependent on $\epsilon$. The results for two levels of noise are shown: $\epsilon_{1}=0.001$ with filled circles, and $\epsilon_{2}=0.1$ with crosses. In case of negligible noise, the points are located along the x-axis (see circles). For larger $\epsilon$, the eigenvalues become more comparable and the plane $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$ fills-up: the data form the shape of a croissant (see crosses). In case of a very large noise (and in case of random fluctuations), the eigenvalues become of the same order, i.e., $\lambda_{2}/\lambda_{1}\sim\lambda_{3}/\lambda_{2}\sim 1$ (not shown). In Section 7, $\epsilon$ is estimated from observations.
@@ 5.1.2 Dipole Alfvén vortex
In Figure 6 column (b), the top panel shows the magnetic field of the dipole vortex. This type of Alfvén vortices is particular, because its axis is inclined with respect to the background magnetic field and it is propagating, see details of the model in appendix A.1.2. The magnetic field components are symmetric in time around the vortex axis (while for the monopole vortex they are anti-symmetric). The magnetic polarisation (third panel from the top) is different for the crossing at the vortex center (black trajectory), and the side crossing (blue trajectory).
Figure 6 (column (b), bottom panel) gives the minimal variance eigenvalues ratios for two noise levels. In case of the low noise, $\epsilon=0.001$, $\lambda_{3}/\lambda_{2}\sim 0$ and $\lambda_{2}/\lambda_{1}\in[0.1,0.6]$ (filled circles). For $\epsilon=0.1$, as for the monopole vortex, both ratios increase: the points in the $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$–plane move towards the upper right corner.